A point mass m moving with velocity u towards the centre of the disc collides with a disc of mass m and radius r resting on a rough horizontal surface. Its collision is perfectly inelastic. Find angular velocity of system after pure rolling starts.


A) 2u/7r      


B) 7u/2r   


C) 5u/2r   


D) 2u/5r

Grade:11

1 Answers

Chetan Mandayam Nayakar
13 years ago

initial linear momentum = mu.  final velocity =v and pure rolloing angular velocity=w.

let total linear impulse = J, total angular impulse = Jr.   linear/angular impulse = change in linear/angular momentum

J=mu-2mv

Jr= mwr2/2 = mvr/2

J=mv/2= mu-2mv,implies v=2u/5,implies w=2u/5r